secA-tanA = K, then sinA = ? sin x + sin y = a ; cos x + cos y = b, then tan(x+y/2) = ? (ans in terms of a and b)

  1.   secA-tanA = K, then sinA = ?
  2.   sin x + sin y = a ; cos x + cos y = b, then tan(x+y/2) = ?  (ans in terms of a and b)

Grade:11

2 Answers

Subhiksha
7 years ago
secA- Tan A =K
1-sinA/ ( 1- sin2A)½ =k
k2= 1- sin A / 1+ sin A
By applying componendo and dividendo,
sinA = 1- k2/1+k2
 
sinx + siny= 2cos(x-y/2)sin(x+y/2) =a
Cosx+ Cosy= 2cos(x-y/2)cos(x+y/2) = b
 
On dividing both the above equations,
= a/b tan(x+y/2)
 
Subhiksha
7 years ago
rewrite the last equationasTan (